OpenStudy (anonymous):

the ratio if john's money to Fred's money is 8:3.If john gives fred $6,john will have twice as much money as fred.How much money does john have before he gives fred money?

terenzreignz (terenzreignz):

This is kind of tricky, you know.

OpenStudy (anonymous):

not for that smart brain of yours

terenzreignz (terenzreignz):

My brain isn't smart. *I* am. ;) HAHA Kidding ^^ What's tricky is explaining it to you. Let me think of a way...

OpenStudy (anonymous):

lol your funny

OpenStudy (anonymous):

i need to hurry cause my mom is rushing me

terenzreignz (terenzreignz):

This isn't something that we can solve in a hurry :> Maybe you should rest now? :D

OpenStudy (anonymous):

i cant cause im really busy and this is due tomorrow morning

OpenStudy (anonymous):

isnt there something we can work out

OpenStudy (anonymous):

can we finish this at 3:30

terenzreignz (terenzreignz):

Well then, your mom's going to have to wait

OpenStudy (anonymous):

lol yeah

OpenStudy (anonymous):

ok lets start

terenzreignz (terenzreignz):

Okay. I'll try and make this quick. But this is not how it should be done, okay? haha

OpenStudy (anonymous):

okay

terenzreignz (terenzreignz):

Suffice it to say, Fred only has 3/8 of John's money. We're looking for how much John originally had.

OpenStudy (anonymous):

yes

terenzreignz (terenzreignz):

So. John's original amount is x Fred's is 3/8 x

OpenStudy (anonymous):

yes

terenzreignz (terenzreignz):

Now if John gives Fred 6, his amount will become x - 6 because he loses 6, right? What happens to Fred's money?

OpenStudy (anonymous):

he will gain 6

terenzreignz (terenzreignz):

so his amount becomes...?

OpenStudy (anonymous):

9

terenzreignz (terenzreignz):

no. Where did you get that? haha His amount was originally 3/8 x so you just add 6 to that

OpenStudy (anonymous):

6 3/8

terenzreignz (terenzreignz):

\[\Large \frac38x + 6\] more like :P

terenzreignz (terenzreignz):

So John's new amount is x - 6 and that thing up there ^ is Fred's new amount Apparently, with these new amounts, John has twice Fred's money. So, in equations, that'd be \[\Large x - 6 = 2\left[\frac38x + 6\right]\]

terenzreignz (terenzreignz):

Now solve for x as you did last time

OpenStudy (anonymous):

thats really confusing like last time lol

terenzreignz (terenzreignz):

It's just the fraction. But actually, you just distribute the 2 to the two terms inside the [] much like you did with the 6 earlier.

OpenStudy (anonymous):

x-6=6/8+12

OpenStudy (anonymous):

is that correct

terenzreignz (terenzreignz):

It is. You can multiply everything by 8 to get rid of that denominator.

terenzreignz (terenzreignz):

And again, you missed a spot. x's don't just disappear XD

terenzreignz (terenzreignz):

\[\Large x - 6 = 2\left[\frac38\color{red}x + 6\right]\]

OpenStudy (anonymous):

i still don't get it can you just tell me what x is

OpenStudy (anonymous):

please friend

OpenStudy (anonymous):

friends are suppose to help friends

terenzreignz (terenzreignz):

I have a name, you know LOL

OpenStudy (anonymous):

please help me terence

OpenStudy (anonymous):

what is x suppose to be

terenzreignz (terenzreignz):

lol that's just an alias I've been using. I'll multiply by 8, you do the rest. You CAN do it, because you did it before lol 8x - 48 = 6x + 96 Now solve it.

OpenStudy (anonymous):

plus 48 to both sides

OpenStudy (anonymous):

???????????

terenzreignz (terenzreignz):

Yup. You get...?

OpenStudy (anonymous):

or minus 96 to both sides

terenzreignz (terenzreignz):

Better to add 48 to both sides.

OpenStudy (anonymous):

8x=6x+144

OpenStudy (anonymous):

??

terenzreignz (terenzreignz):

good. Next?

OpenStudy (anonymous):

-6x to both sides

terenzreignz (terenzreignz):

Okay. And you get...?

OpenStudy (anonymous):

2x=144 which x equals 72 yes

OpenStudy (anonymous):

!!!!!!!!!!!!!!

terenzreignz (terenzreignz):

And that's the answer. See? You CAN do it. You don't need me to tell you the answer haha

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

do you want your reward

terenzreignz (terenzreignz):

new medal? haha sure ^_^